A car dealer advertised a big sale by stretching a string of banners from the top of the building to the edge of the driveway. If the building is 26 m high and the driveway is 43 m from the building, how long is the string of banners?


Answer by Alan3354(69443)

A car dealer advertised a big sale by stretching a string of banners from the top of the building to the edge of the driveway. If the building is 26 m high and the driveway is 43 m from the building, how long is the string of banners?
------------------
s^2 = 26^2 + 43^2
s^2 = 2525
s =~ 50.25 meters